rupees 22 was divided among 100 childrens so that each boy got 25 paise and each girl 15 paise. find the number of girls

Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52795)

.
  x +   y =  100
25x + 15y = 2200


etc.


Answer by addingup(3677)

By what method do you have to solve this? the substitution method is easier here:
boys x; girls y
x + y = 100 so x = 100-y
25x + 15y = 2200 substitute for x
25(100-y)+15y = 2200
2500-25y+15y = 2200
-10y = -300 divide both sides by -10, and remember that -/- = + so you can ignore the negative sign and just divide 300/10 to get your answer of how many girls.